# Impact of the pre-equilibrium phase for the determination of nuclear geometry in high-energy isobar collisions

### v2

@@ III Interplay between free-streaming and nuclear geometry
Fig. 2 shows the same event at the initial time and at the end of the free-streaming. One sees that the energy density distribution becomes more rounded, with a larger average radius $R\equiv\sqrt{\{r^{2}\}}$ as time flows. Therefore, due to free-streaming, $R$ increases while the eccentricities decrease.
This is shown more precisely in Fig. 3. In this figure, the ratio between a given quantity computed after free-streaming and for the initial condition is used. It is a more effective mean of expressing the influence of the prethermalization phase. The following concise notation is used
@@
Here and below, to determine centrality bin edges, all generated events are used, regardless of the (geometrical) case they belong to in Table 2. Such a procedure is different from the actual experiment, where centrality classification is based on the number of charged tracks together with information from Monte Carlo Glauber simulations for each collision system independently Abdallah et al. (2022). Results are in general shown in bins of width 1% in the 0-5% centrality window and then in bins of 5%; finer bins are used to $\rho_{2}$ and $\rho_{3}$, see below. When statistics demands, the 5 most central bins are merged in a single one of 0-5%. Centrality bin edges can be read from Table 3
@@
In Fig. 3, curves end up grouped by how long the produced system free-streamed: for eccentricities, one group of curves corresponds to $\tau_{FS}=0.4$ fm/c, the middle one to $\tau_{FS}=1.0$ fm/c and the last one to the centrality-dependent $\tau_{FS}$, estimated to be approximately between 1.35 fm/c and 1.49 fm/c from Fig. 1. Free-streaming effects are increasingly stronger (meaning, larger deviations from 1) going from $R$, to $\epsilon_{2}$, then $\epsilon_{3}$ and finally $\epsilon_{4}$. Ratios for the 6 different geometrical configurations exhibit extremely similar shapes which mean they are affected fairly similarly by the free-streaming modeling of the prethermalization phase. A conformal expansion during the free-streaming is assumed. Since this model introduces no interactions as time goes by, in the case of a free-streaming velocity smaller than $c$ the system would have a slower expansion, reducing the effects of the prethermalization phase. An interacting model will introduce correlations on the system that will further slow down the eccentricity reductions. In this way, the magnitude of the effects shown here may be understood as an upper bound. For the subsequent hydrodynamic phase, reducing the free-streaming velocity is important for mitigating the artificial increase of the bulk pressure at the start of the expansion da Silva et al. (2023).
@@
Figure 2: The mid-rapidity energy density slice of a ${}^{96}_{40}$r +${}^{96}_{40}$r event in the 20-30% centrality window is shown (left) initially and (right) after free-streaming.
@@
We now concentrate on effects of varying the geometry, for a fixed free-streaming time. Results are shown for a given nuclear configuration (numbered from 1 to 5, with parameters from Table 1), relative to configuration 6 (expected to be more alike a Zr nucleus). For $\tau_{FS}=0.4$ fm/$c$, changes in the nuclear radius and eccentricities compared to their initial values are just a few percent (cf. Fig. 3) and ratios for a given configuration 1-5 with respect to 6 are essentially equal to the ones obtained at initial condition level, so not shown. Unless otherwise stated, we ran 10 million events minimum-bias to obtain the results shown.
@@
Figure 3: Root mean square radius $R$ and eccentricities $\epsilon_{n}$ at the end of free-streaming phase divided by their initial values. Line styles correspond to different free-streaming times. One million minimum-bias events were used to generate this figure.
@@
Figure 4: Left panels: Ratios between ellipticity computed for cases 1 through 5 and case 6, as function of total initial energy and with different free-streaming times. From top to bottom, the calculation is made for mean ellipticity, second and fourth order cumulants. Same line patterns as in Fig. 3 are used, with the addition of solid line to represent quantities computed for the initial conditions. Right panels: ratios between the curves of the left hand panel with free-streaming time and the ones for the initial conditions.
@@
Figure 4 illustrates two effects: the first column represents the influence of distinct nuclear geometries, while the second column emphasizes the effect of free-streaming. The first column of Figure 4 shows ratios for the mean ellipticity employing the different nuclear configurations considered in this work. Configurations 1 and 2 have very similar ellipticity and differ only due to triaxiality $\gamma$ which should not affect this quantity. Configuration 3 shows a higher ellipticity than configuration 2 even though both configurations only differ by the inclusion of an octupole deformation $\beta_{3}$, indicating the non-trivial interplay between $\beta_{2}$ and $\beta_{3}$ discussed e.g. in Jia (2022a). The decrease in ellipticity from configuration 3 to 4 is expected due to the decrease of the quadrupole deformation. Configurations 4 and 5 have similar ellipticity, with ratios close to 1, indicating that this quantity is insensitive to nuclear radius and diffuseness. The differences in geometry are more apparent for the most central collisions.
@@
For the second order cumulant ratios: $\epsilon_{2}\{2\}_{n}/\epsilon_{2}\{2\}_{6}=\langle\epsilon_{2}\rangle_{n}/\langle\epsilon_{2}\rangle_{6}\times\sqrt{1+(\sigma^{2}_{n}/\langle\epsilon_{2}\rangle^{2}_{n})}^{2}\\ /\sqrt{1+(\sigma^{2}_{6}/\langle\epsilon_{2}\rangle^{2}_{6})^{2}}$. We have checked that the quantities $\sigma^{2}_{n}/\langle\epsilon_{2}\rangle^{2}_{n}$ are nearly equal for n=1-6, so it explains the similarity between the upper and lower panels. The ratios of $\epsilon_{2}\{4\}$ are also shown, even though the error bars are large, because it has been argued that the comparison of $v_{2}\{2\}$ and $v_{2}\{4\}$ or $\epsilon_{2}\{2\}$ and $\epsilon_{2}\{4\}$ for isobars might yield information on nuclear structure Jia et al. (2023). The inset illustrates the effect of changing the diffuseness $a$.
@@
Figure 5 shows the triangularity as a function of the total initial energy $E$ for different choices of the free-streaming times. On the left, by comparing the different nuclear configurations it is possible to verify that this quantity is mainly affected by the octupole deformation $\beta_{3}$ (as expected) and the diffuseness. The effect is more prominent in central collisions, where the nucleus overlap more and their nuclear geometry becomes increasingly important. Other details such as the nuclear radius, triaxiality, and quadrupole deformation do not affect this quantity. From the right hand side, it can be seen that free-streaming has larger effects for triangularity than for ellipticity for cases 1 and 2 (divided by that of case 6): there is a larger splitting for the two different free-streaming times, almost 1% for most centralities, which is comparable with the error bars in experimental triangular flow ratios. For cases 3-5, the effect is a little less pronounced. In any case, the modeling and duration of the prethermalization phase may play a role in a very precise determination of the octopole deformation.
@@
Figure 6 shows a higher-order eccentricity, $\varepsilon_{4}$, as function of the total initial energy $E$ for different free-streaming times. When going from case 4 to case 5, for central collisions, the dependence on the diffuseness parameter is 3% for central collisions and larger for non central ones. In line with results obtained for triangularity, a progressively larger splitting between the results with different free-streaming times is observed.
@@
Given that a good mapping is expected between not just $v_{2}$, $v_{3}$ and $\epsilon_{2}$, $\epsilon_{3}$, but also between the event plane angles and initial state symmetry plane angles $\Phi_{2}$, $\Phi_{3}$, it can be used to study the normalized symmetric cumulant NSC(2,3) Gardim et al. (2017),
@@
$\langle\cos 6(\Phi_{2}-\Phi_{3})\rangle=\frac{\langle\epsilon_{2}^{3}\epsilon_{3}^{2}\cos 6(\Phi_{2}-\Phi_{3})\rangle}{\sqrt{\langle\epsilon_{2}^{6}\rangle\langle\epsilon_{3}^{4}\rangle}}\,.$ (7)
@@
Figure 8: Normalized symmetric cumulant $\epsilon NSC(2,3)$ computed from initial conditions for the 6 geometrical cases.
@@
In Fig. (8), for the normalized symmetric cumulant computed directly from the initial conditions, cases 1, 2, 3 are similar and 4 and 5 are fairly equal. Free streaming seems to have a strong effect, and might change once or twice the sign of the ratio, however error bars are large, so more statistics would be necessary to confirm this. The changes of sign can be traced back to slightly different free-streaming time dependencies for each nuclear geometry considered: as the system evolves in time, the individual quantities $\varepsilon NSC(2,3)$ appearing in this ratio becomes less negative, with each of them crossing zero at different energy values. Case 6 only shows a single change of sign (positive to negative) that always happens for values of energy lower than the ones shown in our plots, meanwhile, cases 2 and 3 for instance, present two clear changes of sign (positive to negative and back to positive), with the second change happening in the most central bin considered. The ratio of two-plane correlators, fluctuates around zero and suffer from lack of statistics so is not shown. Another interesting angle correlator for which data have been obtained is $ac2(3)=\langle v_{2}^{2}v_{4}\cos 4(\Psi_{2}-\Psi_{4})\rangle$. Since it cannot be predicted for all centralities with just $\varepsilon_{4}$ Gardim et al. (2012b, 2015b), we leave it to a future study from hydrodynamics. We note however that it has a good potential to discriminate geometry in central collisions Zhao et al. (2023b).
@@
The observables presented so far, except perhaps $\varepsilon NSC(2,3)$, exhibit minimal sensitivity to the effects of free-streaming and triaxiality. One observable known to be sensitive to triaxiality is the Pearson correlator between event anisotropic flow and event mean transverse momentum $[p_{T}]$ in a given centrality bin Bally et al. (2022a); Aad et al. (2023). This correlator has been identified as a strong candidate for assessing the effects on triaxility in isobars Jia (2022b). Previous studies were limited to $n=2$ and utilized estimators that lack proper correlation with mean transverse momentum: Ref. Giacalone et al. (2021c) showed that a good proxy to estimate this type of correlator is to replace the mean momentum by the ratio between total energy and total entropy in a very narrow centrality bin and both $v_{2}\propto\varepsilon_{2}$ and $v_{3}\propto\varepsilon_{3}$ can be considered. Therefore we calculate the following Pearson correlator:
@@
where the averages are taken in centrality bins of 0.25% centrality and $\sigma_{\varepsilon_{n}^{2}}$ and $\sigma_{E/S}$ are the standard deviation of $\varepsilon_{n}$ and $E/S$ in those centrality bins, respectively.
@@
Our results confirm that the effect the triaxiality parameter $\gamma$ is particularly clear when comparing configurations 1 and 2 for $\rho_{2}$ in the top panel of Fig. 9. We also report that the splitting between the results of these two cases increases with the free-streaming time, starting from $\sim 0.3$ at the initial condition level and reaching $\sim 0.5$ when considering a centrality dependent free-streaming duration. Comparing cases 2 and 3, shows that the value of $\beta_{3}$ has a strong effect, lowering the correlation in the $\sim 5\%-13\%$ centrality range. Moreover, one verifies that the free-streaming time plays an interesting role, inducing a single or a double sign change depending on how long the system stays in the prethermalization phase. While we still lack final experimental results for this correlator, such a signature may eventually help to constrain the duration of the pre-hydrodynamic phase in isobar collisions. Next, reducing the quadrupole deformation while keeping the octupole deformation fixed shifts the correlation to values around unity with a much reduced free-streaming time dependence. For the case 5, the result is brought even closer to unity. Using these results together with the ones presented in Fig. 4, 5 indicates that it might be possible to constrain $\beta_{2}$, $\beta_{3}$ and $\gamma$ simultaneously from observations of anisotropic flow and the correlation of elliptical flow with mean transverse momentum.
We have checked that using different estimators for $[p_{T}]$, the effect of the free-streaming time is still substantial.
Moving to $\rho_{3}$, and assuming that the estimator (8) is still a good proxy for the final state observable (as is the case for high-energy Pb+Pb collisions Giacalone et al. (2021c)), one sees that while the dependence on the triaxiality parameter becomes weaker, the effect of free-streaming time is still visible for cases 1-3. In addition, adding an octupole deformation on top of a quadrupole deformation (case 3) shifts the correlation to even larger values in central collisions, providing a clear separation with cases 1 and 2 where only the quadrupole deformation is present. Cases 4, 5 and 6 all have similar $\rho_{3}$, indicating that this observable is independent of diffuseness and distribution radius in the centrality range considered.
@@
Figure 10: Expected value for the distribution of E/S (in a centrality bin) for case n divided by that of case 6 as function of centrality.
Another possibility to determine the geometry is the study of the distribution of event mean transverse momentum in a given centrality bin, $[p_{T}]$. As argued above, a good estimator for $[p_{T}]$ is E/S. It is expected that a good estimator for $\langle\delta[p_{T}]^{p}\rangle/\langle[p_{T}]\rangle^{p}$ would be $\langle(\delta(E/S))^{p}\rangle/\langle E/S\rangle^{p}$ where $\delta O\equiv O-\langle O\rangle$ (as far as we know this has not been tested throughout). The ratio $\langle E/S\rangle_{n}/\langle E/S\rangle_{6}$ as function of centrality is shown in Fig. 10. It is computed directly from the initial conditions (as E/S is conserved) and exhibit a clear ordering: cases n=1-2 have the highest values for the ratio, above 1, and do not depend on the triaxiality. Case 3 is also above 1 but introducing $\beta_{3}$ has decreased the ratio. For case 4, decreasing $\beta_{2}$ increases the ratio. For case 5, increasing $a$ decreases the ratio. However, all these effects are quite small. Fig. 11 is more interesting: the ratios $\left(\langle(\delta(E/S))^{2}\rangle/\langle E/S\rangle^{2}\right)_{n}/\left(\langle(\delta(E/S))^{2}\rangle/\langle E/S\rangle^{2}\right)_{6}$ are shown and larger effects are observed. Cases n=1-2 have the most extreme values for the ratio, below 0.98 and above 1.01, and do not depend on the triaxiality. Case 3 also reaches above 1.01 but introducing $\beta_{3}$ has pushed the ratio towards 1 in general (by about 2%). For case 4, decreasing $\beta_{2}$ decreases the ratio. For case 5, increasing $a$ pushes the ratio very near 1 except for very central collisions. This quantity might be useful to further constrain the value of $\beta_{2}$ given that curves with small values of this parameter are well separated from those with higher values for central collisions. Finally in Fig. 12, $\left(\langle(\delta(E/S))^{3}\rangle/\langle E/S\rangle^{3}\right)_{n}/\left(\langle(\delta(E/S))^{3}\rangle/\langle E/S\rangle^{3}\right)_{6}$ is shown. It has been claimed Jia (2022b) that this quantity could also be used to constrain $\gamma$, but the effect seems very small. Though the error bars are big, some geometrical effects can be seen, for example, introducing $\beta_{3}$ increases the ratio and diminishing $\beta_{2}$ decreases it.
In Jia (2022a), some quantities comparable to those we computed have be obtained as well but with a different estimator for $[p_{T}]$. As a general rule, the effect of geometry is qualitatively similar.

### v4

@@ III Interplay between free-streaming and nuclear geometry
Eccentricities have long been expected to be good estimators for anisotropic flow coefficients [47, 24, 48]. It has been argued that when there is a free-streaming stage and a Laudau matching is done to switch to a hydrodynamic description, an initial momentum anisotropy is created and could also play a role [49, 50]. More recent papers indicate that for nuclear collisions (and some pre-equilibrium stage), initial spatial eccentricity is a good predictor for elliptic flow (see e.g. [51, 52]). Therefore in this paper, we concentrate on spatial eccentricities.
Fig. 3 shows the same event at the initial time and at the end of the free-streaming (computed for configuration (iii)). One sees that the energy density distribution becomes more rounded, with a larger average radius $R\equiv\sqrt{\{r^{2}\}}$ as time flows. Therefore, due to free-streaming, $R$ increases while the eccentricities decrease.
This is shown more precisely in Fig. 4. In this figure, the ratio between a given quantity computed after free-streaming and for the initial condition is used. It is a more effective mean of expressing the influence of the pre-thermalization stage. The following concise notation is used
@@
Here and below, to determine centrality bin edges, all generated events are used, regardless of the (geometrical) configuration they belong to in Table 2. Such a procedure is different from the actual experiment, where centrality classification is based on the number of charged tracks together with information from Monte Carlo Glauber simulations for each collision system independently [2]. Results are in general shown in bins of width 1% in the 0–5% centrality window and then in bins of 5%; finer bins are used to $\rho_{2}$ and $\rho_{3}$, see below. When statistics demands, the 5 most central bins are merged in a single one of 0–5%. Centrality bin edges can be read from Table 3
Table 3: Definition of the centrality bins as a function of E.
@@
In Fig. 4, curves end up grouped by how long the produced system free-streamed: for eccentricities, one group of curves corresponds to $\tau_{FS}=0.4$ fm/c, the middle one to $\tau_{FS}=1.0$ fm/c and the last one to the centrality-dependent $\tau_{FS}$, estimated to be approximately between 1.35 fm/c and 1.49 fm/c from Fig. 2. Free-streaming effects are increasingly stronger (meaning, larger deviations from 1) going from $R$, to $\varepsilon_{2}$, then $\varepsilon_{3}$ and finally $\varepsilon_{4}$. Ratios for the 6 different geometrical configurations exhibit extremely similar shapes (which mean they are affected fairly similarly by the free-streaming modeling of the pre-thermalization stage) so results are only shown for configurations 2 and 6. In the approach we use, the free-streaming velocity is equal to $c$. In case it is smaller, the system would have a slower expansion, reducing the effects of the pre-thermalization stage. An interacting model will introduce correlations on the system that will also attenuate the effects of the pre-thermalization stage. In this way, the magnitude of the effects shown here may be understood as an upper bound. For the subsequent hydrodynamic phase, reducing the free-streaming velocity is important for mitigating the artificial increase of the bulk pressure at the start of the expansion [53].
@@
Figure 3: The mid-rapidity energy density slice of a ${}^{96}_{40}$r +${}^{96}_{40}$r event in the 20–30% centrality window is shown (left) initially and (right) after free-streaming.
@@
We now concentrate on effects of varying the geometry, for a fixed free-streaming time. Results are shown for a given nuclear configuration (numbered from 1 to 5, with parameters from Table 1), relative to configuration 6 (expected to be more alike a Zr nucleus). For $\tau_{FS}=0.4$ fm/$c$, changes in the nuclear radius and eccentricities compared to their initial values are just a few percent (cf. Fig. 4) and ratios for a given configuration 1–5 with respect to 6 are essentially equal to the ones obtained at initial condition level, so not shown. Unless otherwise stated, we ran 10 million events minimum-bias for each nuclear configuration present in Table 1 to obtain the results shown.
@@
Figure 4: Root mean square radius $R$ and eccentricities $\varepsilon_{n}$ at the end of free-streaming stage divided by their initial values. Line styles correspond to different free-streaming times. One million minimum-bias events were used to generate this figure. Results for all configurations lie close so only those for configurations 2 and 6 are shown.
@@
Figure 5: Left panels: Ratios between ellipticity computed for configurations 1 through 5 and configuration 6, as function of total initial energy and with different free-streaming times. From top to bottom, the calculation is made for mean ellipticity, second and fourth order cumulants. Same line patterns as in Fig. 4 are used, with the addition of solid line to represent quantities computed for the initial conditions. Right panels: ratios between the curves of the left hand panel with free-streaming time and the ones for the initial conditions.
@@
Figure 5 illustrates two effects: the first column represents the influence of distinct nuclear geometries, while the second column emphasizes the effect of free-streaming. The first column of Figure 5 shows ratios for the mean ellipticity employing the different nuclear configurations considered in this work. Configurations 1 and 2 have very similar ellipticity and differ only due to triaxiality $\gamma$ which should not affect this quantity. Configuration 3 shows a higher ellipticity than configuration 2 even though both configurations only differ by the inclusion of an octupole deformation $\beta_{3}$, indicating the non-trivial interplay between $\beta_{2}$ and $\beta_{3}$ discussed e.g. in [15]. The decrease in ellipticity from configuration 3 to 4 is expected due to the decrease of the quadrupole deformation. Configurations 4 and 5 have similar ellipticity, with ratios close to 1, indicating that this quantity is insensitive to nuclear radius and diffuseness. The differences in geometry are more apparent for the most central collisions.
@@
The ratio $\langle\varepsilon_{2}\rangle_{i}/\langle\varepsilon_{2}\rangle_{6}$ is very close to the ratio $\varepsilon_{2}\{2\}_{i}/\varepsilon_{2}\{2\}_{6}$ so only the latter is shown. This is because $\varepsilon_{2}\{2\}_{i}/\varepsilon_{2}\{2\}_{6}=\langle\varepsilon_{2}\rangle_{i}/\langle\varepsilon_{2}\rangle_{6}\times\sqrt{1+(\sigma^{2}_{i}/\langle\varepsilon_{2}\rangle^{2}_{i})}^{2}/\sqrt{1+(\sigma^{2}_{6}/\langle\varepsilon_{2}\rangle^{2}_{6})^{2}}$ and we have checked that the quantities $\sigma^{2}_{n}/\langle\varepsilon_{2}\rangle^{2}_{n}$ are nearly equal for $i=1-6$. The ratios of $\varepsilon_{2}\{4\}$ are also shown, even though the error bars are large, because it has been argued that the comparison of $v_{2}\{2\}$ and $v_{2}\{4\}$ or $\varepsilon_{2}\{2\}$ and $\varepsilon_{2}\{4\}$ for isobars might yield information on nuclear structure [54]. The inset illustrates the effect of changing the diffuseness $a$.
@@
Figure 6 shows the triangularity as a function of the total initial energy $E$ for different choices of the free-streaming times. On the left, by comparing the different nuclear configurations it is possible to verify that this quantity is mainly affected by the octupole deformation $\beta_{3}$ (as expected) and the diffuseness. The effect is more prominent in central collisions, where the nucleus overlap more and their nuclear geometry becomes increasingly important. Other details such as the nuclear radius, triaxiality, and quadrupole deformation do not affect this quantity. From the right hand side, it can be seen that free-streaming has larger effects for triangularity than for ellipticity for configurations 1 and 2 (divided by that of configuration 6): there is a larger splitting for the two different free-streaming times, almost 1% for most centralities, which is comparable with the error bars in experimental triangular flow ratios. For configurations 3–5, the effect is a little less pronounced. In any case, the modeling and duration of the pre-thermalization stage may play a role in a very precise determination of the octopole deformation.
@@
Figure 7 shows a higher-order eccentricity, $\varepsilon_{4}$, as function of the total initial energy $E$ for different free-streaming times. When going from configuration 4 to configuration 5, for central collisions, the dependence on the diffuseness parameter is 3% for central collisions and larger for non central ones. In line with results obtained for triangularity, a progressively larger splitting between the results with different free-streaming times is observed.
@@
Given that a good mapping is expected between not just $v_{2}$, $v_{3}$ and $\varepsilon_{2}$, $\varepsilon_{3}$, but also between the event plane angles and initial state symmetry plane angles $\Phi_{2}$, $\Phi_{3}$, it can be used to study the normalized symmetric cumulant $\varepsilon NSC(2,3)$ [55],
@@
$\langle\cos 6(\Phi_{2}-\Phi_{3})\rangle=\frac{\langle\varepsilon_{2}^{3}\varepsilon_{3}^{2}\cos 6(\Phi_{2}-\Phi_{3})\rangle}{\sqrt{\langle\varepsilon_{2}^{6}\rangle\langle\varepsilon_{3}^{4}\rangle}}\,.$ (7)
@@
Figure 9: Ratio of normalized symmetric cumulant $\varepsilon NSC(2,3)$ computed for the geometrical configurations $i=1-5$ compared to $i=6$.
@@
In Fig. 9, for the normalized symmetric cumulant computed directly from the initial conditions, configurations 1, 2, 3 are similar and 4 and 5 are fairly equal and are shown separately for clarity. Free streaming seems to have a strong effect, and might change once or twice the sign of the ratio, however error bars are large, so more statistics would be necessary to confirm this. There are oscillations for $E<200$ GeV so results are shown only for larger values. The changes of sign can be traced back to slightly different free-streaming time dependencies for each nuclear geometry considered: as the system evolves in time, the individual quantities $\varepsilon NSC(2,3)$ appearing in this ratio becomes less negative, with each of them crossing zero at different energy values. Configuration 6 only shows a single change of sign (positive to negative) that always happens for values of energy lower than the ones shown in our plots, meanwhile, configurations 2 and 3 for instance, present two clear changes of sign (positive to negative and back to positive), with the second change happening in the most central bin considered. At this point, it is interesting to discuss what would happen if $\alpha$ were negative in the definition of $\tau_{FS}$ in case (iii). In the leftmost bin shown in the figure, according to fig. 2, the values of $\tau_{FS}$ are fairly equal so the ratio of $\varepsilon NSC(2,3)$ is the same for $\alpha$ positive or negative. In the rightmost bin, again from fig. 2, $\tau_{FS}$ would decrease from $\approx 1.48$ fm/c to $\approx 1.44$ fm/c. Therefore the value of the ratio of $\varepsilon NSC(2,3)$ would be between the value for variable $\tau_{FS}$ and that for $\tau_{FS}=1$ fm/c, but nearest to the former. So the sign changes (top panel) would not be much affected.
The ratio of two-plane correlators, fluctuates around zero and suffer from lack of statistics so is not shown. Another interesting angle correlator for which data have been obtained is $ac2(3)=\langle v_{2}^{2}v_{4}\cos 4(\Psi_{2}-\Psi_{4})\rangle$. Since it cannot be predicted for all centralities with just $\varepsilon_{4}$ [47, 24], we leave it to a future study from hydrodynamics. We note however that it has a good potential to discriminate geometry in central collisions [56].
@@
The observables presented so far, except perhaps $\varepsilon NSC(2,3)$, exhibit minimal sensitivity to the effects of free-streaming and triaxiality. One observable known to be sensitive to triaxiality is the Pearson correlator between event anisotropic flow and event mean transverse momentum $[p_{T}]$ in a given centrality bin [21, 8]. This correlator has been identified as a strong candidate for assessing the effects on triaxility in isobars [16]. Previous studies were limited to $n=2$ and utilized estimators that lack proper correlation with mean transverse momentum: Ref. [52] showed that a good proxy to estimate this type of correlator is to replace the mean momentum by the ratio between total energy and total entropy in a very narrow centrality bin and both $v_{2}\propto\varepsilon_{2}$ and $v_{3}\propto\varepsilon_{3}$ can be considered. Therefore we calculate the following Pearson correlator:
@@
where the averages are taken in centrality bins of 0.25% centrality and $\sigma_{\varepsilon_{n}^{2}}$ and $\sigma_{E/S}$ are the standard deviation of $\varepsilon_{n}$ and $E/S$ in those centrality bins, respectively. $S$ is computed as $S=\int d^{2}x_{\perp}s\left(N\left[\frac{T_{A}(x,y)^{p}+T_{B}^{p}(x,y)}{2}\right]^{1/p}\right)$ where the entropy density $s$ is obtained from the equation of state.
@@
Our results confirm that the effect the triaxiality parameter $\gamma$ is particularly clear when comparing configurations 1 and 2 for $\rho_{2}$ in the top panel of Fig. 10. We also report that the splitting between the results of these two configurations increases with the free-streaming time, starting from $\approx 0.3$ at the initial condition level and reaching $\approx 0.5$ when considering a centrality dependent free-streaming duration. Comparing configurations 2 and 3, shows that the value of $\beta_{3}$ has a strong effect, lowering the correlation in the $\approx 5\%-13\%$ centrality range. Moreover, one verifies that the free-streaming time plays an interesting role, inducing a single or a double sign change depending on how long the system stays in the pre-thermalization stage. While we still lack final experimental results for this correlator, such a signature may eventually help to constrain the duration of the pre-hydrodynamic phase in isobar collisions. Next, reducing the quadrupole deformation while keeping the octupole deformation fixed shifts the correlation to values around unity with a much reduced free-streaming time dependence. For the configuration 5, the result is brought even closer to unity. Using these results together with the ones presented in Fig. 5, 6 indicates that it might be possible to constrain $\beta_{2}$, $\beta_{3}$ and $\gamma$ simultaneously from observations of anisotropic flow and the correlation of elliptical flow with mean transverse momentum.
We have checked that using different estimators for $[p_{T}]$ (taken from  [52]), the effect of the free-streaming time is still substantial.
Moving to $\rho_{3}$, and assuming that the estimator (8) is still a good proxy for the final state observable (as is the case for high-energy Pb+Pb collisions [52]), one sees that while the dependence on the triaxiality parameter becomes weaker, the effect of free-streaming time is still visible for configurations 1–3. In addition, adding an octupole deformation on top of a quadrupole deformation (configuration 3) shifts the correlation to even larger values in central collisions, providing a clear separation with configurations 1 and 2 where only the quadrupole deformation is present. Configurations 4, 5 and 6 all have similar $\rho_{3}$, indicating that this observable is independent of diffuseness and distribution radius in the centrality range considered.
@@
Figure 11: Expected value for the distribution of $E/S$ (in a centrality bin) for configuration $i$ divided by that of configuration 6 as function of centrality.
The $\alpha<0$ case can be studied in a similar way as done for Fig. 9, one sees that it would not alter the general behavior of $\rho_{2}$ and $\rho_{3}$.
Another possibility to determine the geometry is the study of the distribution of event mean transverse momentum in a given centrality bin, $[p_{T}]$. As argued above, a good estimator for $[p_{T}]$ is $E/S$. It is expected that a good estimator for $\langle\delta[p_{T}]^{p}\rangle/\langle[p_{T}]\rangle^{p}$ would be $\langle(\delta(E/S))^{p}\rangle/\langle E/S\rangle^{p}$ where $\delta O\equiv O-\langle O\rangle$ (as far as we know this has not been tested throughout). The ratio $\langle E/S\rangle_{i}/\langle E/S\rangle_{6}$ as function of centrality is shown in Fig. 11. It is computed directly from the initial conditions (as E/S is conserved) and exhibit a clear ordering: configurations $i=1-2$ have the highest values for the ratio, above 1, and do not depend on the triaxiality. Configuration 3 is also above 1 but introducing $\beta_{3}$ has decreased the ratio. For configuration 4, decreasing $\beta_{2}$ increases the ratio. For configuration 5, increasing $a$ decreases the ratio. However, all these effects are quite small. Fig. 12 is more interesting: the ratios $\left[\langle(\delta(E/S))^{2}\rangle/\langle E/S\rangle^{2}\right]_{i}/\left[\langle(\delta(E/S))^{2}\rangle/\langle E/S\rangle^{2}\right]_{6}$ are shown and larger effects are observed. Configurations $i=1-2$ have the most extreme values for the ratio, below 0.98 and above 1.01, and do not depend on the triaxiality. Configuration 3 also reaches above 1.01 but introducing $\beta_{3}$ has pushed the ratio towards 1 in general (by about 2%). For configuration 4, decreasing $\beta_{2}$ decreases the ratio. For configuration 5, increasing $a$ pushes the ratio very near 1 except for very central collisions. This quantity might be useful to further constrain the value of $\beta_{2}$ given that curves with small values of this parameter are well separated from those with higher values for central collisions. Finally in Fig. 13, $\left[\langle(\delta(E/S))^{3}\rangle/\langle E/S\rangle^{3}\right]_{i}/\left[\langle(\delta(E/S))^{3}\rangle/\langle E/S\rangle^{3}\right]_{6}$ is shown. It has been claimed [16] that this quantity could also be used to constrain $\gamma$, but the effect seems very small. Though the error bars are big, some geometrical effects can be seen, for example, introducing $\beta_{3}$ increases the ratio and diminishing $\beta_{2}$ decreases it.
In [16], some quantities comparable to those we computed have be obtained as well but with a less efficient estimator for $[p_{T}]$. As a general rule, the effect of geometry is qualitatively similar.